what is (2x+3y) (2x+y) {foil it} the back of the book says its 4x^2+8xy+3y^2 but idk why plz explain :)
the lesson is multiplying binomials

Answer by scott8148(6628)

FOIL ___ First terms, Outside terms, Inside terms, Last terms

F ___ 2x * 2x = 4x^2

O ___ 2x * y = 2xy

I ___ 3y * 2x = 6xy

L ___ 3y * y = 3y^2

combining like terms ___ 4x^2 + 8xy + 3y^2